OpenStudy (anonymous):

Jill rides her horse 5 miles due North and then 3 miles due East. Rounded to the nearest tenth what is the magnitude and direction of Jills ride?

OpenStudy (anonymous):

@shifuyanli

OpenStudy (anonymous):

5.8 mi; 31° east of north

OpenStudy (anonymous):

distance is AB. use pythag method to calculate AB i.e. AB = 52+32−−−−−−√=34−−√=5.8309518948453004708741528775456 rounded to 5.8 use trig function to get that angle. opposie/adjacent = tan(x) i.e. tan(x) = 3/5 = 0.6 x = arctan(3/5) = 30.9637565321 rounded to 31 Hence tha above answer. Let me know if that is correct.

OpenStudy (anonymous):

@HackberryAbby18 ???

OpenStudy (anonymous):

Yes I believe so

OpenStudy (anonymous):

okay (: